40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 5218     Entry time: Sat Aug 13 01:52:07 2011
Author: Yoichi 
Type: Update 
Category: LSC 
Subject: Feed forward delay 
Yoichi, Koji

While I was testing the feed forward cancellation, I noticed that the
cancellation was not perfect.
The test I did was the following.
I injected the same signal to both DARM and MICH feedback filters.
This was done by injecting a signal into the excitation point of
the ASDC PD, then changing the input matrix elements so that the signal
goes to both DARM and MICH.
Then in the FFC, MICH signal was fed forward to DARM by the gain of -1.
Ideally, this should completely eliminate the DARM FB signal.
In reality, it did not.

The first PDF compares the spectrum of the injected noise (white noise,
red curve) with the spectrum of the signal after the FFC (blue curve).
At higher frequencies, the cancellation becomes poor.
It suggests that this is caused by some delay in the FFC.
I also took a transfer function from the injection point to the signal
after the FFC (second attachment).
I fitted the measured TF with a theoretical formula of
where dt is the time delay and f is the frequency.
The fitting is very good, and I got dt = 0.8msec ~ 13 samples for 16kHz.
13 samples is something very large.

The cause of the delay was suspected to be the shared memory communication
between different processes.
I moved all the FFC blocks to c1lsc.mdl.
Then the cancellation becomes perfect. The signal after the FFC is
completely zero, so I couldn't even make a TF measurement.

This results suggest that a large delay of 13 samples is induced
when you use shared memory to send signals round trip.
We should make simpler models, just passing signals back and forth
via shared memory, dolphin network or GE FANAC RFM to check the
delays more precisely.

For the moment, the FCC is included in the c1lsc model.
The MEDM screens were modified to account for this change.
c1ffc is stopped and removed from rtsystab.
Attachment 1: Spe1.pdf  24 kB  | Hide | Hide all
Attachment 2: TFFitting.png  224 kB  | Hide | Hide all
ELOG V3.1.3-